### v1

# Compositional Hardness of Code in Large Language Models - A Probabilistic Perspective

Abstract
A common practice in large language model (LLM) usage for complex analytical tasks such as code generation, is to sample a solution for the entire task within the model’s context window. Previous works have shown that subtask decomposition within the model’s context (chain of thought), is beneficial for solving such tasks. In this work, we point a limitation of LLMs’ ability to perform several sub-tasks within the same context window – an in-context hardness of composition, pointing to an advantage for distributing a decomposed problem in a multi-agent system of LLMs. The hardness of composition is quantified by a generation complexity metric, i.e., the number of LLM generations required to sample at least one correct solution. We find a gap between the generation complexity of solving a compositional problem within the same context relative to distributing it among multiple agents, that increases exponentially with the solution’s length. We prove our results theoretically and demonstrate them empirically.
1 Introduction
Large language models (LLMs), based on the transformer archietecture, Vaswani et al. (2017), have become very efficient problem solvers in many domains, such as broad-scoped question answering, writing assistance, teaching, and more Brown (2020); Radford et al. (2019); OpenAI (2023); Bubeck et al. (2023); Nori et al. (2023); West (2023). Yet their analytical skills, such as coding capabilities, are slow to develop - Chen et al. (2021b); Li et al. (2022a); Alp (2023); Ridnik et al. (2024) show that even with millions of generations, LLMs may not produce a single correct solution to competitive coding problems. Dziri et al. (2024) show that LLMs solve compositional tasks such as long multiplication and dynamic programming without developing systematic problem-solving skills.
One way to empower LLMs in analytical tasks, is to use subtask decomposition, otherwise known as chain of thought (COT) - a method in which an LLM breaks down a problem to smaller, more manageable tasks, solves them, and integrates it into a solution. The method has been empirically demonstrated by Wei et al. (2022), that show reasoning capabilities of language models improve when they are prompted to break down a task. Its efficiency has also been studied theoretically - Wies et al. (2022); Malach (2023) prove that through the autoregressive nature of language models, problems that cannot be solved directly, can be solved by subtask decomposition, Merrill & Sabharwal (2023) prove that while transformers are limited in the computational problems they can solve directly, using a polynomial number of intermediate steps, they can represent any polynomial time Turing Machine, and Sanford et al. (2024) provide a similar result on expressing more complex arithmetic circuits using more steps of COT.
Yet even with task decomposition, there is a limitation to transformer based models on analytical tasks with COT, due to their limited ability to compose functions - Sanford et al. (2024) show single layer attention can only learn pairwise relations between tokens, limiting the ability to integrate intermediate steps for a fixed model size, and the work of Peng et al. (2024) shows that iterative composition over a domain is limited by the size of the model, even when COT is used. Thus while theoretically possible, some tasks require an arbitrarily long COT for an LLM to solve. However, in practice, LLMs are limited in their context length - beyond the constraint of context length during training, Hsieh et al. (2024) introduce the RULER benchmark, for measuring the utility of LLMs on long context tasks. They show that in practice many models can perform tasks only on a much shorter context length than they were trained on. Similarly, Ebrahimi et al. (2024) show empirically that LLMs are limited in random access to tokens within the context, in the bit parity problem. Consequently, even though COT can in theory allow an LLM to solve arbitrarily complex analytical problems, in practice, they will be limited by the effective context length.
A rising approach to remedy this limitation is to solve problems through the use of multi-agent systems, that tackle complex problems through the use of agents, where each agent is an LLM instance that solves a different aspect of the problem. While it has been used for simulating social interactions, Park et al. (2023); Li et al. (2023); Pang et al. (2024), it has also been shown as an effective tool for analytical problem solving. This can be done by decomposing a large task and distributing the sub-tasks between agents. Ishibashi & Nishimura (2024) use this method for building large code bases and Liu et al. (2023), use a dynamic LLM-agent network for solving code problems and analytical tasks.
In this work, we theoretically study a compositional hardness of coding problems originating from context processing limitations of LLMs, and the resultant effectiveness of a multi-agent system over a single model instance in composite coding problems. We model a composite coding problem using a pair of simpler coding problems, where the output to the composite problem can be a basic manipulation on the outputs to the pair of problems (e.g. composition, product of outputs, concatenation, logical function of the outputs, etc.), such that the solution to the problem can be obtained from concatenating the solutions to the problem pair. The model’s usefulness on a coding task, is quantified by a generation complexity metric (definition 2) - the number of LLM generations required to sample at least one correct solution. The appeal of this metric, is that due to the existence of code testing units, it suffices to turn an LLM into a good program candidate generator and simply output a candidate that is correct Kambhampati et al. ; Thakur et al. (2023); Luo et al. (2024). The single LLM instance solves the entire problem, while the multi-agent system is comprised of two agents, each is an LLM instance tasked with solving one of the problems in the pair. Thus the LLM’s helpfulness in the single instance case is the generation complexity for the composite problem, while in the multi-agent case, it is the product of generation complexities of the pairs of problems, as a correct solution is attained when independently sampling a correct solution to each problem.
We theoretically model an LLM as an autoregressive model, where a solution is sampled token by token, based on the hidden representations. When combining two problems whose solutions are grammatically similar but semantically different (such as different coding problems), we assume this combination injects noise into the model’s representations during solution generation for each sub-problem (assumption 1), which we denote as screening. We show that a compositional problem may have an exponential gap in generation complexity relative to the product of sub problems’ generation complexities (theorem 1), meaning an exponential hardness of composition in-context. Essentially, the model is less capable of solving two problems if they are presented within the same context, than if presented in separate contexts. This points to an advantage of decomposing a problem not only within the same LLM context, but to distribute the problems among multiple agents (i.e. solve each sub-task within a different context). Additionally, this result provides a view of the model’s effective context length through the lens of screening, which is the model’s ability to isolate the relevant context at each decoding step – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length. We validate our assumptions and results experimentally on Llama 3 by constructing composite code problems.
2 Framework
2.1 Generation Complexity
We focus on coding problems, meaning each problem is written in natural language and is solved by a function, and the goal is for the model to generate a code that implements it.
Definition 1.
Let $L$ be a programming language. Let $x$ be a natural language description of a problem, that is solved by function $f$, then a computer program $y$ is a correct solution to $x$, if it implements $f$.
We formally define the generation complexity of a program to a problem as the inverse success rate of a model to generate a correct solution to the problem:
Definition 2.
For a problem $x$ and a natural language distribution $\mathbb{P}$ over $\Sigma^{*}$, the generation complexity of $x$ w.r.t. $\mathbb{P}$, is:
$N(\mathbb{P},x)=\frac{1}{\sum_{y\in\textit{correct solutions}}\mathbb{P}(y|x)}$ (1)
Intuitively, the generation complexity is the number of program candidates needed to be sampled from the distribution to get a program that solves the problem.
In this work, we consider simple compositional problems, which are decomposable to two smaller problems, and their concatenated solutions solve the compositional problem. This captures cases such as a composition of two code functions, or simple manipulations on outputs of code functions (e.g. concatenation, product, etc.). For a compositional problem $x$, decomposable to $x_{1},x_{2}$, we quantify the in-context hardness of composition as the ratio of generation complexity to the full problem $N(\mathbb{P},x)$ with the product of generation complexities for the sub-problems $N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$. If the latter is much smaller, this points to an advantage for distributing the sub-problems between different instances of an LLM (multiple agents), as it would require fewer generations to sample a correct solution. This defines a compositional hardness of coding problems that originates from context processing limitations, as the compositional problem can be mathematically equivalent to solving both problems, yet seeing them both in the same context reduces the model’s performance on them.
2.2 Screening in Autoregressive Models
Here we introduce a source of hardness in code composition based on the autoregressive nature of LLMs. Typically, latent representations of the model contain information about the context beyond the next token prediction, e.g. the structure of the solution to problems Ye et al. (2024). Thus when composing two code problems, we expect the representations during the generation of the second program to contain information about the first program and vice versa, which is grammatically similar (same programming language) but semantically very different. As a result, this additional information creates noise that can harm the generation process.
Formally, we denote by $r^{(L)}(x)$ the model’s last hidden layer representation of the prompt $x$, by $U$, the model’s unembedding matrix (hidden dimension to vocabulary). The logit of the $i$’th token is thus defined as the token’s score before the softmax operation: $\langle r^{(L)}(x),U^{T}e_{i}\rangle$, where $e_{i}$ is the one-hot vector of the token. The probability distribution at each decoding step is the softmax applied to the logits, $\mathbb{P}_{LLM}(i|x)=softmax(\langle r^{(L)}(x),U^{T}e_{i}\rangle)$.
In the process of generating a solution to a compositional code problem, $x$, that is implicitly or explicitly decomposable to $x_{1}$ and $x_{2}$, the model will implement a solution to the first part $y_{1}$ and then to the second part $y_{2}$. The sequence is generated based on the hidden representations. Informally, we expect the representation of the solution to the first problem, $y_{1}$, within the compositional problem, $x$, to be a noisy version of the solution’s representation in the non-compositional problem, $x_{1}$:
$r^{(L)}(x\oplus y_{1})=r^{(L)}(x_{1}\oplus y_{1})+noise$ (2)
Similarly, the representation of the second problem’s solution, $y_{2}$, in the compositional problem, $x$, is expected to be a noisy version of the solution’s representation in the non-compositional problem, $x_{2}$:
$r^{(L)}(x\oplus y_{1}\oplus y_{2})=r^{(L)}(x_{2}\oplus y_{2})+noise$ (3)
Essentially, this means the model attempts to generate the same solutions as in the non-compositional case, but noise may interfere in the process. The projection of this noise onto the dictionary creates noise in the logits during decoding, which can lead the model to make mistakes.
As the two problems $x_{1},x_{2}$ and their solutions $y_{1},y_{2}$ may be different semantically, we do not expect the noise to "push" the model towards the correct solutions more than to incorrect solutions, thus when projected onto the vocabulary, $V$, the noise on the logit of the correct token minus the noise on an incorrect token, $\langle noise,U^{T}e_{i_{\textit{correct next token}}}\rangle-\langle noise,U^{T}e_{i_{\textit{an incorrect next token}}}\rangle$, should be symmetric on average. Additionally it should be bounded within some range $[-M,+M]$, as it changes the hidden representation to a finite extent. In practice, we only expect this to be true for the high probability tokens, as the vocabulary $V$ is very large, and some low probability tokens may be systematically enhanced. To avoid this issue, we make our assumptions only on the weighted average of the noise, where the weights are given by the probability mass that the model assigns them. This way, low probability tokens with asymmetric noise or large norms receive low weight and are averaged with the noise of other tokens. Denote by $\mathbb{P}(i|context)$ the probability assigned to the $i$’th token given the context. We will make our assumptions on the weighted average of the noise on the incorrect token logits minus correct token logits:
$X=\frac{\sum_{i\in V/\{\textit{correct next token}\}}\mathbb{P}(i|context)\langle noise,U^{T}e_{i}-U^{T}e_{i_{\textit{correct next token}}}\rangle}{\sum_{i\in V/\{\textit{correct next token}\}}\mathbb{P}(i|context)}$ (4)
Assumption 1.
Denote by $X$ the weighted noise on the logits as defined in equation 4. We assume that at every given decoding step it is a continuous, symmetric random variable and bounded within $[-M,+M]$ for some $M>0$.
In experiment subsection 4.3 we show the noise satisfies these assumptions, with $M\approx 3-4$.
2.3 Effect of noise on decoding
While the noise onto the logits, $X$, averages to zero, its effect on the decoding process does not. The probability of each token in a decoding step is changed to:
$\mathbb{P}(i|context)\rightarrow\mathbb{P}^{\prime}(i|context)\leq\frac{\mathbb{P}(i|context)}{\mathbb{P}(i|context)+(1-\mathbb{P}(i|context))e^{X}}$ (5)
The denominator, $\mathbb{P}(i|context)+(1-\mathbb{P}(i|context))e^{X}$ can be thought of as a renormalizing term, which redistributes the probability of the tokens. For $X=0$, the token’s probability does not change, for $X<0$ it increases and for $X>0$ it decreases. Note that if $\mathbb{P}(i|context)\in(\epsilon,1-\epsilon)$ for $\epsilon>0$ (the model has finite confidence), then on average, the noise decreases the probability $\mathbb{P}(i|context)$, by a factor of $exp(-\Delta(\epsilon,X))$, where $\Delta$ is the renormalizing term’s mean:
$\Delta(\epsilon,X):=\mathbb{E}_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ (6)
Intuitively $\Delta$ is the average renormalization of the correct token’s probability. In experiment subsection 4.3, we calculate $\Delta(\epsilon,X)$ empirically as a function of $\epsilon$, and find that for $\epsilon=0.1$ for example, $\Delta\approx 0.2$. The consequence of this, is that on average, most long sequences have their probability reduced by the noise, while few random long sequences have their probability greatly enlarged. Since for long coding problems most sequences are incorrect, the probability of a correct solution getting enlarged is small. We formally show this in the next section. To do so, we will use concentration inequalities, for which we note that:
$|\log(\epsilon+(1-\epsilon)e^{X})|<M$ (7)
Thus the renormalizing term’s variance is also bounded:
$\sigma^{2}(\epsilon,X):=Var_{X}[\log(\epsilon+(1-\epsilon)e^{X})]\leq M^{2}$ (8)
3 Results
Here we show that composing two coding problems can be significantly harder than solving each on its own. A natural quantification of compositional hardness is the gap between generation complexity of the problem’s components and the complete problem. For example, we say that composition is hard if:
$N(\mathbb{P},x)\gg N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ (9)
While it is easy if:
$N(\mathbb{P},x)\approx N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ (10)
The rational behind this, is that $N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ is the number of attempts required to independently sample a correct solution to $x_{1}$ and to $x_{2}$, while $N(\mathbb{P},x)$ is the number of attempts required to sample a solution to problem that integrates both problems. In an easy composition scenario, the model solves the sub-problems to the best of its abilities as it would if each was solved independently. In a hard composition scenario, seeing both problems combined reduces its performance on each sub-problem, and it is better to use a multi-agent system, by feeding the model the subproblems in different contexts and sample solutions independently.
As there are typically more incorrect solutions to coding problems than correct ones, the random noise inserted into the logits generally harms the model’s performance. The following lemma quantifies an exponential decrease in the model’s probability of a correct solution to a compositional problem, $\mathbb{P}(y_{1}\oplus y_{2}|x)$, relative to the probabilties of the sub-problem solutions $\mathbb{P}(y_{1}|x_{1})\cdot\mathbb{P}(y_{2}|x_{2})$:
Lemma 1.
Let $\epsilon,\delta\in(0,1)$, and $M>0$. Let $x$ be a compositional problem and $y_{1}\oplus y_{2}$ a solution, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Suppose that the noise injected to the logits as defined in equation 4, satisfies assumption 1 - for all decoding steps, it is continuous, symmetric and bounded within $[-M,+M]$. Suppose further that the probability assigned to the correct token at each decoding step is bounded within $[\epsilon,1-\epsilon]$. Denote by $\Delta:=\mathbb{E}_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ and $\sigma^{2}:=Var_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ the renormalizing term’s mean and variance (as defined in equations 6 and 8 respectively). Under the assumption, $\Delta,\sigma$ are strictly positive, and if $|y_{1}|+|y_{2}|>\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\ln\frac{1}{\delta}$, where $h(x)=(x+1)\ln(1+x)-x>0$, we have with probability of at least $1-\delta$ that:
$\mathbb{P}(y_{1}\oplus y_{2}|x)\leq\mathbb{P}(y_{1}|x_{1})\cdot\mathbb{P}(y_{2}|x_{2})e^{-\frac{\Delta\cdot(|y_{1}|+|y_{2}|)}{4}}$ (11)
Where $\mathbb{P}(y|x)$ is the probability assigned to $y$ by the model, given context $x$.
The proof is presented in appendix A. The intuition behind this result is that as there are typically more incorrect choices to make when generating code, random noise usually reduces the probability for sequences with finite confidence. Thus most sequences get their probability reduced, while very few random sequences get a large increase, and these are usually not correct solutions.
The assumption on bounded probability for the correct token $[\epsilon,1-\epsilon]$ implies we are considering solutions where the model has high but limited confidence in each decoding step. While LLMs are often very confident, in practice, when sampling, nucleus sampling is commonly used, where sampling only occurs if the model is not overly confident. e.g. with $p=0.95$, if a token’s probability $P$, is larger than $0.95$ or much smaller than $0.05$, then the probability is rounded to 00 or $1$. Thus when considering probabilities of sequences, it suffices to look only at decoding steps where the model is not too confident, hence we can consider $\epsilon\approx 0.05$.
In subsection 4.3, we see empirically that for $\epsilon=0.1$, $\Delta\approx 0.2,\sigma\approx 1.5,M\approx 4$, making $\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\approx 200$, thus for solutions with length $>200$ tokens, the results apply.
In practice, there may be multiple solutions to the same problem, e.g. multiple implementations of the same function. So it is necessary to take all of them into account when considering the generation complexity. With this taken into account using a union bound, we obtain the following result of an exponential gap in generation complexity between a composition of problems and the sub-problems, indicating a compositional hardness that is exponential in the solution’s length:
Theorem 1.
Let $\epsilon,\delta\in(0,1)$, and $N,M>0$. Let $x$ be a compositional problem, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Denote by $L_{1},L_{2}$ the minimal solution length to $x_{1},x_{2}$ respectively, and the total number of solutions to $x$ by $N$. Define $\Delta,\sigma$, the renormalizing term’s mean and variance (as defined in equations 6 and 8 respectively) and by $M$ the bound on the logit noise (assumption 1). Under the assumptions of lemma 1, they are strictly positive, $\Delta,\sigma,M>0$, and if the minimal solution length $L_{1}+L_{2}$, satisfies $L_{1}+L_{2}>\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\ln\frac{N}{\delta}$, where $h(x)=(x+1)\ln(1+x)-x>0$, we have with probability of at least $1-\delta$ that the generation complexity (definition 2) satisfies:
$N(\mathbb{P},x)\geq N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})\cdot e^{\frac{\Delta\cdot(L_{1}+L_{2})}{4}}$ (12)
The proof is presented in appendix B. We see that longer problems become harder to solve due to the noise injected into the decoding steps by previously generated tokens. This shows that a model that fully utilizes its context in decoding (i.e. the next token probability distribution is explicitly a function of all the previous tokens), can have a hard time mixing different concepts due to the screening effect. This result implies that for long coding problems, it is more beneficial to distribute sub-tasks between different instances of the LLM, and not expose it to the full context. Additionally, this result provides a view of the model’s effective context length through the lens of screening – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length. The intuition to the result is that there are more incorrect solutions than correct ones, so the random noise usually reduces their total probability mass.
We note that for the union bound of theorem 1 to hold, we need the number of solutions to be bounded by an exponential in the length of the solution, $L_{1}+L_{2}$:
$N<\delta\cdot exp\left((L_{1}+L_{2})\cdot\frac{\sigma^{2}}{M^{2}}h\left(\frac{3\Delta\cdot M}{4\sigma^{2}}\right)\right)$ (13)
Typically, we expect the number of solutions to a problem to grow exponentially with the length of the shortest solution, $y$:
$N\sim exp(c\cdot|y|)$ (14)
As each line of code can be implemented slightly differently, lines may be interchanged, different variable names, etc. Still, the exponential’s coefficient is expected to be small as most sequences are not correct solutions to the problem due to the constraints between the tokens (e.g., once a variable name is chosen, it is fixed throughout the solution). With the empirical values calculated in subsection 4.3, we see empirically that the coefficient of the exponential number of solutions is $\approx 0.005$, thus for a solution of length $L_{1}+L_{2}=1,600$ tokens, we have $N<\delta\cdot exp(8)=3,000\cdot\delta$ solutions.
4 Experiments
In this subsection we test the assumptions and results of our theoretical part. We create simple compositional coding problems with pairs of problems from the humaneval benchmark Chen et al. (2021a) and code contests dataset Li et al. (2022b). First, we show the results of theorem 1, stating that composition is typically harder than independently solving the subproblems. Then, we show explicit indications for the exponential length dependence of compositional hardness, by comparing probabilities of solutions with/without composition, as theoretically suggested in lemma 1. Finally, we look at our assumption 1 on the noise inserted into the logits of the second problem, and observe that it is indeed large enough to interfere with the decoding process (assumption 1). The experiments were performed on Llama-3-8B-Instruct Dubey et al. (2024).
4.1 Generation Complexity Results
Here we test the actual generation complexity of an LLM to different problems corresponding to scenarios from our theoretical results. As proposed in the theoretical section, an LLM may be able to solve two problems with low generation complexity, but their composition might take a significantly larger generation complexity if the model was not explicitly trained on such a task. To test this, we built a set of composite problems based on the Human Eval benchmark Chen et al. (2021a) and code contests dataset Li et al. (2022b).
To create composite problems, we took pairs of problems from Human Eval and code contests, and created from each a problem whose solution requires to explicitly solve both problems. Additionally, for harder problems, we created compositions of problems from code contests dataset. The three templates used were:
  • (Human Eval) Problem 1 has a True/False output, Problem 2 has an arbitrary output. Composition is to solve problem 1, then if its output is true, print the second problems’ output, otherwise, print "-1":
  • (Human Eval) Problem 1 and Problem 2 have an integer output. Composition is to solve both and print the product of their outputs.
  • (Code Contests) Problem 1 and problem 2 are problems from the dataset. Composition is to read the inputs sequentially and print the outputs sequentially.
We denote the composite problem of $x_{1}$ and $x_{2}$ as $x_{1}\oplus x_{2}$. For each problem we sampled 200 solutions and evaluated the generation complexity as the inverse of percentage of correct solutions.
Figure 1 shows the cumulative distribution function (CDF) of the compositional hardness - the value along the $y$ axis is the CDF of $\frac{N(\mathbb{P},x)}{N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})}$, i.e.  percentage of compositional problems, in which $\frac{N(\mathbb{P},x)}{N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})}$ is smaller than a given value. For example, with $a=5$, the corresponding value on the $y$ axis, is the percentage of problems in which composition requires up to ×$\times 5$ more generations relative to solving independently. As can be seen, for most of the problems, composition generation complexity is larger than the product of generation complexities of the components (up to factors of $10-20$).
Figure 1: Cumulative distribution function for the ratio of generation complexity using composition, $N(\mathbb{P},x)$, to product of generation complexities for the standalone problems, $N(\mathbb{P},x_{1})\cdot N(\mathbb{P},x_{2})$ (corresponding to the multi-agent generation complexity). The x axis denotes values for the ratio of generation numbers required to solve the problem in the two cases (composition vs multi-agent), the y axis is the percentage of problems in which the ratio is no larger than this value (e.g. for $a=5$, the y axis value is the percentage of problems where less than ×$\times 5$ samples are required in composition than multi-agent). (a) For the human eval composition. As can be seen in most of the cases, composition requires twice more samples, and for some problems 10 times more samples. (b) For the code contests composition. As can be seen the majority of problems have a factor of at least 5, and some up to 20.
As seen in theorem 1, we have:
$N(\mathbb{P},x_{1}\oplus x_{2})\gg N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})$ (15)
Furthermore, we see that in code contests, where the problems typically have longer solutions (few hundreds of tokens), and the problems are harder thus the model is less certain, the compositional hardness is much larger.
Assuming one only has access to end-to-end verifier of the composite problem, it is more advantageous to generate independently for each problem then to generate for the composite problem, giving the advantage to a multi-agent system over a single LLM instance.
4.2 Exponential Length Dependence of Compositional Hardness
Here, we used the same compositions as in the previous subsection, and measured the difference in probabilities of correct solutions with vs without composition as a function of the number of tokens in the solution. The "correct" solutions were taken as the canonical solutions of the dataset, as they are "neutral" in the sense that solutions generated by the model with/without composition may have different styles, and measuring the probability of these generated sequences may create an artificial bias in favor of one of the two.
As suggested by lemma 1, we expect to see on average:
$\log\frac{\mathbb{P}(y_{1}|x_{1})}{\mathbb{P}(y_{1}|x)}\geq\frac{\Delta}{4}\cdot|y_{1}|\textit{ ~{}~{}~{},~{}~{}~{} }\log\frac{\mathbb{P}(y_{2}|x_{2})}{\mathbb{P}(y_{2}|x\oplus y_{1})}\geq\frac{\Delta}{4}\cdot|y_{2}|$ (16)
Meaning that the log ratio of correct solutions without/with composition increases linearly with length. As can be seen in figure 2, an exponential increase in probability of the correct solution is observed without composition relative to with composition, as a function of length of the code. As the y axis is plotted in the log domain, the linear curve approximates $\Delta/4$, which takes a of $\approx 0.05$. This means a decrease of $e^{-1}$ in success rate for every 20 tokens. This matches the typical increase in generation complexity seen in figure 1. In practice, $\Delta$ is likely larger for some compositions, as the average is over many sequences, yet still has high fluctuations.
Figure 2: Ratio of correct solution probability with vs without composition. An exponential trend is observed as a function of length.
4.3 Assumptions
Here we look at our assumption of noise inserted into the logits (assumption 1).
Noise distribution:
Using the same compositions as before, we measure the change to the logits of correct tokens with vs without composition. As in the theoretical assumption, we subtract this with the mean change in the logits of the incorrect tokens:
$X=\frac{\sum_{i\in incorrect}\mathbb{P}_{i}\langle noise,U^{T}e_{i}-U^{T}e_{correct}\rangle}{\sum_{i\in incorrect}\mathbb{P}_{i}}$ (17)
We calculate this over different sequences. As can be seen in figure 3, the change in logits creates a noise that is symmetric, bounded, $X<M\approx 4$, and has finite absolute deviation $E[|X|]>0$, in accordance with assumption 1.
Figure 3: Change in logits of correct tokens minus incorrect tokens due to composition.
Estimation of $\Delta(\epsilon,X)$:
To estimate $\Delta(\epsilon,X)$ we approximate the noise as a Gaussian with mean 0 and try two values standard deviation $\sigma=1$ and $\sigma=2$, then calculate $\Delta(\epsilon,X)\approx mean_{\{X_{i}\}}[\epsilon+(1-\epsilon)e^{X}]$. The values are plotted in figure 4, match the estimation of $\Delta\approx 0.05-0.2$ from the above subsection.
Figure 4: Numerical Estimation of $\Delta(\epsilon,X)$.
Estimation of $\sigma(\epsilon,X)$:
Similarly, we estimate $\sigma(\epsilon,X)$. The values are plotted in figure 5, showing typical values of $\sigma\approx 1-2$.
Figure 5: Numerical Estimation of $\sigma(\epsilon,X)$.
5 Discussion
In this work, we point a limitation of LLMs’ ability to perform several sub-tasks within the same context window – an in-context hardness of composition, pointing to an advantage for distributing a decomposed problem in a multi-agent system of LLMs over using a single LLM instance. The hardness of composition is quantified by the generation complexity metric, i.e., the number of LLM generations required to sample at least one correct solution. We found an exponential gap between generation complexity of solving a composition problem within the same context relative to distributing it among multiple agents. This is attributed to the transformer’s nature of using all tokens in the context simultaneously for decoding, which inserts noise into generated sequences, caused from mixing in sub-tasks’ latent representations (screening). Consequently, even if a model has the ability to perform two tasks, it may not be able to perform them both within the same context. This provides a view of the model’s effective context length through the lens of screening, which is the model’s ability to isolate the relevant context at each decoding step – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length.
Acknowledgments
We thank Yoav Levine and Noam Wies for fruitful discussions. This research was supported by the ERC (European Research Council) and the ISF (Israel Science Foundation).
Appendix A Proof of Lemma 1
We start by proving the following lemma:
Lemma 2.
Let $p\in(0,1)$ and $X$ a random variable over real numbers. Denote by: $\Delta\left(p,X\right)=E\left[\log{\left(p+\left(1-p\right)e^{X}\right)}\right]$. Claim: For a continuous symmetric random variable, X, we have $\Delta\left(p,X\right)>0$. Furthermore, if $p\in(\epsilon,1-\epsilon)$, then $\Delta(p,X)>\Delta(\epsilon,X)=\Delta(1-\epsilon,X)$
Proof:
Denote by $\rho(X)$ the density function of $X$, then:
$E\left[\log{\left(p+\left(1-p\right)e^{X}\right)}\right]=\int_{-\infty}^{0}{\log{\left(p+\left(1-p\right)e^{x}\right)}\rho(x)}+\int_{0}^{\infty}{\log{\left(p+\left(1-p\right)e^{x}\right)}\rho(x)}=$ (18)
Switching sign of the integration variable in the second term:
$=\int_{0}^{\infty}{\log{\left(p+\left(1-p\right)e^{-x}\right)}\rho(-x)}+\int_{0}^{\infty}{\log{\left(p+\left(1-p\right)e^{x}\right)}\rho(x)}=$ (19)
Using the symmetry condition on $X$:
$=\int_{0}^{\infty}{\log{\left(p+\left(1-p\right)e^{-x}\right)}\rho(x)}+\int_{0}^{\infty}{\log{\left(p+\left(1-p\right)e^{x}\right)}\rho(x)}=$ (20)
$=\int_{0}^{\infty}{\left(\log{\left(p+\left(1-p\right)e^{-x}\right)}+\log{\left(p+\left(1-p\right)e^{x}\right)}\right)\rho(x)}=$ (21)
Again, applying the symmetry of $X$:
$=\frac{1}{2}\int_{-\infty}^{\infty}\left(\log{\left(p+\left(1-p\right)e^{-x}\right)}+\log{\left(p+\left(1-p\right)e^{x}\right)}\right)\rho\left(x\right)$ (22)
The integrand is positive:
$\mathrm{log}(p+(1-p)e^{x})+\mathrm{log}(p+(1-p)e^{-x})\mathrm{=}$ (23)
$=\log{(p^{2}+\left(1-p\right)^{2}+p\left(1-p\right)\left(e^{x}+e^{-x}\right)}=$ (24)
$\log{(1+}p\left(1-p\right)\left(e^{x}+e^{-x}-2\right))$ (25)
For $x=0$, the argument inside the log equals $1$, otherwise, the argument inside the log is always larger than $1$, since $e^{x}+e^{-x}-2>0$. Hence the integrand is positive everywhere except in $x=0$, meaning the integral is positive for any continuous symmetric $X$. Thus, for any $p$ the expectation value is positive.
$E\left[\log{\left(p+\left(1-p\right)e^{X}\right)}\right]=\Delta\left(p,X\right)>0$ (26)
Furthermore it is symmetric around $p=1/2$, and monotonic for $p\rightarrow 0$ and $p\rightarrow 1$:
$\mathrm{log}(p+(1-p)e^{x})+\mathrm{log}(p+(1-p)e^{-x})\mathrm{=}$ (27)
$=\log{(p^{2}+\left(1-p\right)^{2}+p\left(1-p\right)\left(e^{x}+e^{-x}\right)}=$ (28)
$\log{(1+}p\left(1-p\right)\left(e^{x}+e^{-x}-2\right))$ (29)
As can be seen, it is monotonic with $p(1-p)$, and if $p\in(\epsilon,1-\epsilon)$, it is minimal for $p=\epsilon,1-\epsilon$.
Proof of Main Lemma:
We now move to the proof of the main lemma. We show the exponential growth with the length of the solution to the second problem, $y_{2}$ (the proof for the exponential dependence on $y_{1}$ is identical). Let us look at the probability that the model assigns the $n$’th token in the sequence $y_{2}$ in the compositional problem. It is given by the softmax on the projections of the final hidden layer representation on the vocabulary given the context:
$P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)=\frac{e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{y_{2}\left[n\right]}\right\rangle}}{e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{y_{2}\left[n\right]}\right\rangle}+\Sigma_{i\in\left[V\right]}e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{i}\right\rangle}}$ (30)
According to assumption 1, $\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{i}\right\rangle=\left\langle r^{(L)}\left(x_{2}\oplus y_{2}\left[:n\right]\right),U^{T}e_{i}\right\rangle+X_{i}$, meaning the model is receiving a noisy version of the representation to the problem it is trying to solve, where $X_{i}$ is the noise onto the $i$’th token.
$=\frac{P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)e^{X_{y_{2}\left[n\right]}}}{P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)e^{X_{y_{2}\left[n\right]}}+\Sigma_{i\in\left[V\right]}P\left(i\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)e^{X_{i}}}$ (31)
For brevity, denote $P_{0}=P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)$, and $P_{i}=P\left(i\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)$.
$=\frac{P_{0}e^{X_{0}}}{P_{0}e^{X_{0}}+\Sigma_{i\in\left[V\right]/{0}}P_{i}e^{X_{i}}}$ (32)
Now, using the AMGM inequality:
$\leq\frac{P_{0}e^{X_{0}}}{P_{0}e^{X_{0}}+\left(1-P_{0}\right)e^{\Sigma_{i\in\left[V\right]}\frac{P_{i}X_{i}}{1-P_{0}}}}=\frac{P_{0}}{P_{0}+\left(1-P_{0}\right)e^{\Sigma_{i\in\left[V\right]}\frac{P_{i}X_{i}}{1-P_{0}}-X_{0}}}$ (33)
Now, denote $X=\Sigma_{i\in\left[V\right]}\frac{P_{i}(X_{i}-X_{0})}{1-P_{0}}$, according to assumption 1, it is a symmetric, continuous random variable, bounded between $[-M,+M]$. Rewriting the above as:
$=P_{0}e^{\log(P_{0}+(1-P_{0})e^{X})}$ (34)
So the correct token probability with composition $P_{0}^{\prime}$ is decreased the factor in the exponent $\rightarrow P_{0}e^{\log(P_{0}+(1-P_{0})e^{X})}$, relative to the probability without composition, $P_{0}$. For a full sequence, we apply the probability chain rule and obtain the following:
$P\left(y_{2}\middle|x\oplus y_{1}\right)=P\left(y_{2}\right|x_{2})e^{-\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}^{i}+\left(1-P_{0}^{i}\right)e^{X_{i}}\right)}}$ (35)
Where $P_{0}^{i}$ is the probability of the correct token in the $i$’th step without composition. Now, because $E_{X_{i}}\left[\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}\right]=\Delta\left(P_{0},X_{i}\right)>0$, from the above lemma. We get a sum of random variables with mean that is larger than zero. We will use a concentration inequality to bound it.
We start with bounding the random variables:
$\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}<\log{\left(P_{0}+\left(1-P_{0}\right)e^{M}\right)}\leq M$ (36)
$\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}>\log{\left(P_{0}+\left(1-P_{0}\right)e^{-M}\right)}\geq-M$ (37)
Also notice that since $P_{0}\in(\epsilon,1-\epsilon)$, the above lemma implies:
$\Delta\left(P_{0},X_{i}\right)>\Delta\left(\epsilon,X_{i}\right)$ (38)
Thus from linearity of the expectation value:
$E\left[\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}\right]>\left|y_{2}\right|\cdot\Delta(\epsilon,X)$ (39)
We can then apply Bennet’s inequality:
$P\left(\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}>\frac{1}{4}\left|y_{2}\right|\cdot\Delta(\epsilon,X)\right)\geq=1-\exp{\left(-\frac{L\sigma^{2}}{M^{2}}h{\left(\frac{3\Delta M}{4\sigma^{2}}\right)}\right)}$ (40)
Where $\sigma^{2}$ is the variance of each summand, $\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}$ (which is positive, since the variable is continuous), $M$ is its bound, $\Delta=\Delta(\epsilon,X)$ is a lower bound for the expectation value, and $h(x)=(1+x)\log(1+x)-x$. Thus for:
$\left|y_{2}\right|>\frac{M^{2}}{\sigma^{2}h(\frac{3\Delta M}{4\sigma^{2}})}\cdot\ln{\frac{1}{\delta}}$ (41)
We obtain the with probability of at least $1-\delta$ that:
$P\left(y_{2}\middle|x\oplus y_{1}\right)<P\left(y_{2}\right|x_{2})e^{-\frac{\Delta}{4}|y_{2}|}$ (42)
Using the same idea for $y_{1}$, we obtain:
$P\left(y_{1}\middle|x\right)<P\left(y_{1}\right|x_{1})e^{-\frac{\Delta}{4}|y_{1}|}$ (43)
Thus together, if $\left|y_{1}\right|+\left|y_{2}\right|>\frac{M^{2}}{\sigma^{2}h(\frac{3\Delta M}{4\sigma^{2}})}\cdot\ln{\frac{1}{\delta}}$, we have with probability $1-\delta$ that:
$P\left(y_{1}\oplus y_{2}\middle|x\right)<P\left(y_{1}\right|x_{1})P\left(y_{2}\right|x_{2})e^{-\frac{\Delta}{4}(|y_{1}|+|y_{2}|)}$ (44)
Appendix B Proof of theorem 1
Now, suppose there are $N$ solutions to the problem, then we need to use a union bound. We get the result of the lemma 1 over all sequences with probability:
$\left(1-\exp{\left(-\frac{L\sigma^{2}}{M^{2}}h{\left(\frac{3\Delta M}{4\sigma^{2}}\right)}\right)}\right)^{N}\geq 1-N\exp{\left(-\frac{L\sigma^{2}}{M^{2}}h{\left(\frac{3\Delta M}{4\sigma^{2}}\right)}\right)}$ (45)
Require this to equal:
$=1-\delta$ (46)
Thus for:
$\left|y_{1}\right|+\left|y_{2}\right|>\frac{M^{2}}{\sigma^{2}h(\frac{3\Delta M}{4\sigma^{2}})}\cdot\ln{\frac{N}{\delta}}$ (47)
We obtain with probability of at least $1-\delta$ that all solutions satisfy the result of lemma 1. If the minimal solution length is $L_{1}+L_{2}$, for all solutions to satisfy the inequality, we require:
$L_{1}+L_{2}>\frac{M^{2}}{\sigma^{2}h(\frac{3\Delta M}{4\sigma^{2}})}\cdot\ln{\frac{N}{\delta}}$ (48)
Thus we have with probability $1-\delta$ that:
$N(\mathbb{P},x)=\frac{1}{\sum_{y_{1},y_{2}\in\textit{correct solutions}}\mathbb{P}(y_{1}\oplus y_{2}|x)}\geq~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (49)
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{1}{\sum_{y_{1},y_{2}\in\textit{correctsolutions}}\mathbb{P}(y_{1}|x_{1})\mathbb{P}(y_{2}|x_{2})}e^{\frac{\Delta}{4}(L_{1}+L_{2})}=$ (50)
$=\frac{1}{\sum_{y_{1}\in\textit{correct solutions for $x_{1}$}}\mathbb{P}(y_{1}|x_{1})\sum_{y_{1}\in\textit{correct solutions for $x_{2}$}}\mathbb{P}(y_{2}|x_{2})}e^{\frac{\Delta}{4}(L_{1}+L_{2})}$ (51)
$=N(\mathbb{P},x_{1})N(\mathbb{P},x_{2})e^{\frac{\Delta}{4}(L_{1}+L_{2})}$ (52)